salam sir please help me by providing the answer of below question.
how many triangles can be formed by joining 12 points ,7 of which lies on the same straight line.
1
Expert's answer
2011-12-01T11:30:13-0500
We can make C(12,3) different combinations of these points, but C(7,3) of them will not create triangle. So, there can be formed C(12,3) - C(7,3) = 220 - 35 = 185 triangles.